a 500kg is into a vertical firing.If the exhaust speed of gas is 800m/s to attain 20m/s upward acceleration and overcomes gravitational pull,.the amount of gas ejected per second?

Grade:11

1 Answers

Puneeth
10 years ago
Since it overcome the gravitational pull,and since there is only one external force acting on the system therefore f(ext)=0.We also know that f(ext)+f(th)=Ma,where f(th) is the upthrust produced .therefore applying the values we get 500*20=0+dm/dt(800)=dm/dt=10000/800=12.5 kg/s.